the burr distribution has a probability function for a continuous random variable x given by f(x)={(ckxc-1)/(1+xc)k+1 x>o 0 otherwise} show that the E(x) = kB(k+1/c,1/c+1)


1
Expert's answer
2021-11-29T03:44:38-0500

beta function:

"B(x,y)=\intop^1_0t^{x-1}(1-t)^{y-1}dt"


"E(x) =\\int^{\\infin}_0xf(x)dx=\\int^{\\infin}_0x\ufeff\ufeff\\frac{ckx^{\ufeff\ufeff\ufeffc-1}}{(1+x^c)^{k+1}}dx"


let "u=(1+x^c)^{-1}\implies x=(\frac{1-u}{u})^{1/c}"

"du=-cx^{c-1}(1+x^c)^{-2}dx=-cx^{c-1}u^2dx"

then:


"E(x) =k\int^{0}_1(\frac{1-u}{u})^{1/c}u^{k+1}(-u^{-2})du=k\int^{1}_0u^{(k-1/c)-1}(u-1)^{(k+1/c)-1}du="


"=kB(k+1\/c,1\/c+1)"